OpenStudy (anonymous):

HOW DO you solve sin ^2-tanxcos^ 2=0 on the interval

OpenStudy (anonymous):

(sin(x))^2-tan(x)(cos(x))^2=0

OpenStudy (anonymous):

tan(x)=sin(x)/cos(x) sin(x)[sin(x)-cos(x)]=0

OpenStudy (anonymous):

===> sin(x)=0 or sin(x)=cos(x) sin(x)=0 x=asin(0)=0 --- sin(x)=cos(x)=> x=45 deg